OpenStudy (anonymous):

the position of a dragonfly that is flying parallel to the ground is given as a function of time by r=[2.90m+(0.0900m/s^2)t^2]i-(0.0150m/s^3)t^3j. (a) At what value of t does the velocity vector of the insect make an angle of 30.0 clockwise from the +x-axis? (b) At the time calculated in part (a), what are the magnitude and direction of the acceleration vector of the insect?

OpenStudy (anonymous):

Position vector: r(t) = (2.9 + 0.09t^2) i - 0.015t^3 j Velocity vector is derivative of position: v(t) = dr/dt = 0.18t i - 0.045t^2 j a.) We're looking for the angle of the velocity vector, which we can find using our trig identity: tan(theta) = opposite/adjacent For our v(t) we get tan(theta) = -0.045t^2 / 0.18t = -0.25t theta(t) = arctan(-0.25t) And we want to know at what time t does theta(t) = -30 degrees: arctan(-0.25t) = -30 -0.25t = tan(-30) t = tan(-30)/-0.025 t = 2.31s b.) Acceleration vector is derivative of velocity vector: a(t) = dv/dt = 0.18 i - 0.09t j At t = 2.31s the acceleration vector is 0.18 i - (0.09)(2.31) j = 0.18 i - 0.208 j In polar coordinates that's 0.275m/s^2 @-49.1 degrees.

OpenStudy (anonymous):

@dmancine thank you for your answer,but when I put the answer in masterphysics "-49.1 degrees" was + not negative.

OpenStudy (anonymous):

I can't see how r(t), v(t), or a(t) would have a positive j component at t>0. Can you see where I might have gone wrong?

OpenStudy (anonymous):

actually i didnt understand the part you found 49,1 degrees. how did you find it?

OpenStudy (anonymous):

I converted the rectangular coordinate vector 0.18i - 0.208j into polar coordinates. Do you understand how to do that?